Can you please help me find the solution to this problem?
-3x(x-7)(-4+11)=0
Thank you.

Found 3 solutions by stanbon, cthom56, JonathanM:
Answer by stanbon(75887)

find the solution to this problem?
-3x(x-7)(-4+11)=0
---------------
if x=0 or is x=7 the product on the left side will equal zero
===============
Cheers,
Stan H.

Answer by cthom56(2)

-3x(x-7)(-4+11)=0
-3(7)(7-7)(-4+11)
-21(0)(7)=
0x7=0

Answer by JonathanM(3)

- 3x ( x - 7 ) ( -4 + 11 ) = 0

(- 3x^2 + 21x ) ( 7 ) = 0

-21x^2 + 147x = 0

-7x ( 3x - 21 )= 0 take the value of x,

x = 0 ; x = 7

Cheers; Jon Po